# RMAAT: Astrocyte-Inspired Memory Compression and Replay for Efficient Long-Context Transformers

### v1

@@ Appendix D AMRB Algorithm
This appendix provides the detailed pseudo-code for the Astrocytic Memory Replay Backpropagation (AMRB) algorithm, as described in Section 3.4. AMRB enables memory-efficient training of the recurrent RMAAT architecture.
Input: $m_{1}$: Initial memory state (input to step $t=1$).
Output: $m_{T+1}$: Updated memory state (output after step $t=T$).
1
// Store initial input state $m_{1}$
2
/* Forward Pass */
3 for $t=1$ to $T$ do
  $m^{\prime}_{t+1}\leftarrow\text{Model}(x_{t},m_{t})$ ;
  // Compute intermediate state (no grad)
  // Apply retention factor
4  if $t<T$ then
     Append $m_{t+1}$ to replay_bufferreplay\_buffer ;
     // Store input state $m_{t+1}$ for next step’s recomputation
5
6
/* Backward Pass */
Initialize $\nabla m_{T+1}\leftarrow 0$ ;
// Init gradient for state after last step
7 for $t=T$ to 11 do
  Retrieve $m_{t}$ from replay_bufferreplay\_buffer ;
  // Get input state for segment $t$
  $m^{\prime}_{t+1},o_{t}\leftarrow\text{Model}(x_{t},m_{t})$ ;
  // Recompute segment $t$ (track grads)
  Compute loss $L_{t}\leftarrow\text{loss\_function}(o_{t})$ ;
  // Compute loss for current step
  Perform backpropagation: $L_{t}.\text{backward()}$ ;
  // Compute param grads $\partial L_{t}/\partial\theta_{t}$, etc.
  // Compute $\nabla m_{t}$ via chain rule
8
9Save $m_{T+1}$ for the next rollout’s update
Algorithm 1 Astrocytic Memory Replay Backpropagation (AMRB)
@@
  • The indexing convention here follows t=1Tt=1...T, where $m_{t}$ is the input memory state to segment $t$, and $m_{t+1}$ is the output state.
@@ E.1 Hyperparameters
Key hyperparameters for RMAAT across the evaluated LRA tasks are summarized in Table S2. Consistent settings were used for iso-architecture baselines where applicable, with task-specific adjustments primarily for sequence length handling (Number of Segments) and training schedule (Epochs, Learning Rate). All models were trained from scratch using the AdamW optimizer and CrossEntropyLoss where applicable.
Table S2: Key Hyperparameters for RMAAT on LRA Tasks.
@@ Appendix F Additional Results
Table S3: RMAAT Accuracy (%) Sensitivity on ListOps ($8K$).
@@
Table S4: RMAAT Accuracy (%) Sensitivity on Text ($4K$).

### v2

@@ Appendix D AMRB Algorithm
This appendix provides additional explanatory notes on the Astrocytic Memory Replay Backpropagation (AMRB) algorithm presented in Section 3.4 (Algorithm 1), focusing on gradient flow and implementation details.
@@
  • The indexing convention in the algorithm follows t=1Tt=1...T, where $m_{t}$ is the input memory state to segment $t$, and $m_{t+1}$ is the output state.
@@ E.1 Hyperparameters
Key hyperparameters for RMAAT across the evaluated LRA tasks are summarized in Table 4. Consistent settings were used for iso-architecture baselines where applicable, with task-specific adjustments primarily for sequence length handling (Number of Segments) and training schedule (Epochs, Learning Rate). All models were trained from scratch using the AdamW optimizer and CrossEntropyLoss where applicable.
Table 4: Key Hyperparameters for RMAAT on LRA Tasks.
@@ Appendix F Additional Results
Table 5: RMAAT Accuracy (%) Sensitivity on ListOps ($8K$).
@@
Table 6: RMAAT Accuracy (%) Sensitivity on Text ($4K$).